OpenStudy (anonymous):

Solve. x2 + 5x + 6 = 0 A. {2, 3} B. {1, –6} C. {–2, –3} D. {–1, 6}

OpenStudy (anonymous):

Step 1: Factor left side of equation.

Nnesha (nnesha):

easy method to factor quadratic equations AC method or headphone method find two number when you multiply them you should get product of AC and when you add or subtract them you should get middle term \[\huge\rm Ax^2+Bx+C=0\] a=leading coefficient b=middle term c=constant term

OpenStudy (anonymous):

Step 2: Set factors equal to zero

OpenStudy (anonymous):

so i'm between B and C then

Nnesha (nnesha):

why ?

OpenStudy (anonymous):

ahhhhhhhhhhhhhhhhhh

Nnesha (nnesha):

that's not an answer.

Nnesha (nnesha):

it's not that hard just look at there a= leading coefficient which is 1 b=5 =6 what two number you should multiply to get 6 and when you add them you should get 5

OpenStudy (anonymous):

(x+2)(x+3)=0

OpenStudy (anonymous):

x+2=0 or x+3=0 x=−2 or x=−3